# State Permutation Control in Non-Hermitian Multiqubit Systems with Suppressed Non-Adiabatic Transitions

### v1

@@ Acknowledgements.
Appendix C Resolving the problem of exceptional points at infinity for single qubits in Eq. (3)
Here, we briefly address the apparent issue of spectral divergence at EPs in the non-interacting qubits NHH, according to Eq. (8). This problem can be resolved using the framework of complex projective spaces, which provides an elegant method to handle infinities. In particular, what appears as an infinity in a given space may correspond to a well-defined point in a higher-dimensional space.
Such a higher-dimensional $3\times 3$ NHH can be introduced as follows
where $\gamma=\Delta\sin\left({\rm Re}[\phi]\right)$. The first two diagonal elements of the NHH are continuous functions on the plane $(x,y)$. The third element is also a continuous function except the interval $[x=0,|y|\leq 1]$, where it becomes rather a piecewise function, i.e., it is not differentiable but remains finite on that interval.
$\lambda_{1}^{\rm EP}=\lambda_{2}^{\rm EP}=0.$ (28)
Figure 6: This Figure is an extension of Figs. 3(d) and 4(d), i.e., the calculated state transfer fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between evolving state $|\psi(t)\rangle$ and one of the eight initial eigenstates $|\psi_{k}\rangle$ of the three-qubit NHH. The panels in the plot are related to different combinations of the modulated system parameters $J_{mn}$ in Eq. (IV.1), and which thus correspond to permutation generators in Eq. (IV.3). The panels (a),(b),(c) and (d) correspond to the permutation generators $p_{3}$, $p_{4}$, $p_{5}$ and $p_{6}$, respectively. The remaining system parameters are the same as in Figs. 3(d) and 4(d).
Appendix D Calculated fidelity for other state permutations according to Eq. (IV.3)
In Figs. 3(d) and 4(d) we plotted the fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between evolving state $|\psi(t)\rangle$ and one of the eight initial eigenstates $|\psi_{k}\rangle$ of the three-qubit NHH that correspond to the first generators of the permutation group $\cal G$ in Eq. (IV.3). For completeness, here in Fig. 6 we plot the fidelity of the remaining permutation generators, thus confirming their form in Eq. (IV.3).
We also summarize various eigenstate flips produced by the permutation generators $p_{i}$ in Table 1. From this Table it is clear that one can always find such a combination of modulated system parameters in Eq. (IV.1) that enables a state transfer between any two given eigenstates in one or more full cycles $T$.
$\psi_{1}$ $p_{2}$ $p_{4},p_{5}$ $p_{1},p_{3}$ $p_{6}$
$\psi_{2}$ $p_{2}$ $p_{5}$ $p_{1},p_{3},p_{4},p_{6}$
$\psi_{3}$ $p_{4},p_{5}$ $p_{1}$ $p_{2}$ $p_{3},p_{6}$
$\psi_{4}$ $p_{5}$ $p_{1}$ $p_{2},p_{4}$ $p_{3}$
$\psi_{5}$ $p_{1},p_{3}$ $p_{2},p_{4}$ $p_{5}$
$\psi_{6}$ $p_{1},p_{3},p_{4},p_{6}$ $p_{2}$ $p_{5}$
$\psi_{7}$ $p_{3},p_{6}$ $p_{5}$ $p_{1},p_{2},p_{4}$
$\psi_{8}$ $p_{6}$ $p_{3}$ $p_{5}$ $p_{1},p_{2},p_{4}$
@@
Appendix E Effects of perturbations on dynamical state permutations in three-qubit system
@@
However, as demonstrated in Figs. 7 and 8, for small levels of disorder ($\epsilon\approx 0.01$), the desired state permutations remain feasible. By appropriately tuning the system parameters, the state transfer fidelity can still be kept at high rates.
@@
Figure 7: State transfer fidelity for the perturbed three-qubit NHH, according to Eq. (1) and Eq. (29) for a perturbation $\epsilon=0.01$. This figure corresponds to the unperturbed case shown in Fig. 3(d). The upper (lower) row of panels demonstrates the scenario when the dynamical winding is performed in the counterclockwise $+\omega$ (clockwise $-\omega$) direction. The time period is chosen as $T=50$ [arb. units], the remaining system parameters are the same as in Fig. 3.
@@
Figure 8: State transfer fidelity for the perturbed three-qubit NHH, according to Eq. (1) and Eq. (29) for a perturbation $\epsilon=0.01$. This figure corresponds to the unperturbed case shown in Fig. 4(d). The upper (lower) row of the panels shows the scenario when the dynamical winding is performed in the counterclockwise $+\omega$ (clockwise $-\omega$) direction. The time period is chosen as $T=120$ [arb. units], and $g_{0}=1.5$, the remaining system parameters are the same as in Fig. 4.

### v2

@@ Acknowledgements.
Appendix E Realizing the non-Hermitian Hamiltonian in Eq. (1) via the $\beta$-dyne detection method
The NHH of Eq. (1) can also be experimentally implemented in the $\beta$-dyne detection scheme [56]. This method allows introducing complex driving fields in the system’s Hamiltonian. That approach is based on a specific homodyne-detection monitoring scheme of a quantum system surrounded by a Markovian environment into which the system’s excitations are leaking out.
To understand the main idea of the approach, let us first consider a master equation, in the Lindblad form, which defines the dynamics of an open quantum system [82]
$\frac{\partial\hat{\rho}(t)}{\partial t}=-i\left[\hat{H}_{\rm eff},\hat{\rho}(t)\right]+\sum_{\mu}\gamma_{\mu}\hat{a}_{\mu}\hat{\rho}(t)\hat{a}_{\mu}^{\dagger}.$ (37)
In Eq. (37), the Hermitian operator $\hat{\rho}$ is the system’s density matrix, $\hat{a}_{\mu}$ are annihilation bosonic operators playing the role of dissipation operators accounting for quantum jumps, and $\hat{H}_{\rm eff}$ is an effective Hamiltonian consisting of both coherent and incoherent parts, namely,
$\hat{H}_{\rm eff}=\hat{H}-\sum\limits_{\mu}\dfrac{i\gamma_{\mu}}{2}\hat{a}_{\mu}^{\dagger}\hat{a}_{\mu},$ (38)
where $\hat{H}$ describes coherent dynamics.
The effective NHH $\hat{H}_{\rm eff}$ defines the system’s continuous non-unitary evolution, i.e., its dynamics between two successive quantum jumps. This continuous evolution of an open quantum system can be accessed via the postselection procedure [83, 84, 29, 85, 86].
Now one notices that the Lindblad master equation in Eq. (37) is invariant under the following affine transformations [56]
$\hat{K}_{\mu}$ $=$ $\hat{a}_{\mu}+\beta_{\mu}\hat{1}$,
$\hat{H}^{\prime}$ $=$ $\hat{H}-\sum_{\mu}\frac{i\gamma_{\mu}}{2}\left(\beta_{\mu}^{*}\hat{a}_{\mu}-\beta_{\mu}\hat{a}_{\mu}^{\dagger}\right).$ (39)
As a result, the effective NHH $\hat{H}_{\rm eff}$ changes as following
$\hat{H}_{\rm eff}(\beta)$ $=\hat{H}-\sum_{\mu}\frac{i\gamma_{\mu}}{2}\left(\beta_{\mu}^{*}\hat{a}_{\mu}-\beta_{\mu}\hat{a}_{\mu}^{\dagger}\right)$ (40)
[…]
From Eq. (40) it is evident that, via this invariance, the NHH can acquire an additional term, which emulates complex driving fields.
Physically speaking, the affine transformations in Eq. (E) correspond to the scenario where one postselects the leaking photons via homodyne detection. In a typical homodyne detection setup, the leaking photons are mixed with an intense coherent field on a balanced beam splitter, resulting in displaced photons with a large coherent component, $\beta\gg 1$. In the regime where the displacement is weak, i.e., $\beta\ll 1$ however, this detection scheme is known as $\beta$-dyne detection [56]. The $\beta$-dyne method can be implemented by mixing photons emitted from the system with an intense coherent field using an unbalanced beam splitter with low reflectivity. This $\beta$-dyne unraveling method thus allows one to introduce virtual complex driving fields in the system’s effective non-Hermitian Hamiltonian [56].
This approach can be employed for the experimental realization of the NHH in Eq. (1) in an open quantum optical system operating in the single-photon regime. Indeed, in this regime (i.e., truncating the Fock space at the single photon), the Pauli matrices in Eq. (1) can be represented by the annihilation ($\hat{a}$) and creation ($\hat{a}^{\dagger}$) bosonic operators, namely,
$\sigma_{x}^{k}$ $=$ $\hat{a}_{k}+\hat{a}_{k}^{\dagger},$
$\sigma_{y}^{k}$ $=$ $-i(\hat{a}_{k}-\hat{a}_{k}^{\dagger}),$
$\sigma_{z}^{j}\sigma_{z}^{k}$ $=$ $(1-2\hat{n}_{j})(1-2\hat{n}_{k}),$
where the photon number is defined as $\hat{n}_{i}=\hat{a}^{\dagger}_{i}\hat{a}_{i}$.
Now consider a Hermitian Hamiltonian describing a weakly driven three-mode system in a cavity with the Kerr nonlinearity in the single-photon regime
$\hat{H}=\sum\limits_{i=1}^{3}\Delta_{i}\hat{n}_{i}+\sum\limits_{j<k}U_{jk}\hat{n}_{j}\hat{n}_{k}+\sum\limits_{i=1}^{3}(\eta_{i}^{*}\hat{a}_{i}+\eta_{i}\hat{a}_{i}^{\dagger}).$
We further assume that the cavity is lossy; that is, each mode $\hat{a}_{i}$ is characterized by the photon decoherence rate $\gamma_{i}$. Then, by monitoring the leaking photons from the cavity via the $\beta$-dyne detection, one can induce the continuous non-unitary evolution in the system governed by the effective NHH which can be identified with that in the Eq. (1). Namely, by combining Eqs. (40)–(E), and by introducing the following relations
$\Delta_{i}$ $=$ $-2\sum\limits_{j<k}J_{jk}-\dfrac{i\gamma_{i}}{2},$
$U_{jk}$ $=$ $4J_{jk},$
$\gamma_{k}\beta_{k}$ $=$ $2\left[{\rm Im}(B_{x}^{k})+i{\rm Im}(B_{y}^{k})\right],$
$\eta_{k}$ $=$ $(B_{x}^{k})^{*}+i(B_{y}^{k})^{*},$ (43)
the resulting effective $\beta$-dyne NHH can be exactly mapped to the NHH in Eq. (1), provided that $\gamma_{i}=\gamma$ for $\forall i$, such that the modes’ decay rate $\gamma/2$ can be gauged away. However, the latter condition can be eased when the difference $|\gamma_{i}-\gamma_{j}|\neq 0$, $i\neq j$, but it is still small enough ensuring that the protocol remains stable. We also refer the interested reader to Ref. [56], which provides a detailed discussion on the physics of the $\beta$-dyne detection scheme.
@@
Appendix F Effects of perturbations on dynamical state permutations in a three-qubit system
@@
However, as demonstrated in Figs. 11 and 12, for small levels of disorder ($\epsilon\approx 0.001$), the desired state permutations remain feasible. By appropriately tuning the system parameters, the state transfer fidelity can still be kept at high rates.
@@
Figure 11: State transfer fidelity for the perturbed three-qubit NHH, according to Eq. (1) and Eq. (44) for a perturbation $\epsilon=0.001$. This figure corresponds to the unperturbed case shown in Fig. 3(d). The upper (lower) row of panels demonstrates the scenario when the dynamical winding is performed in the counterclockwise $\omega$ (clockwise $\omega$) direction. The time period is chosen as $T=500$ [arb. units], the remaining system parameters are the same as in Fig. 3.
@@
Figure 12: State transfer fidelity for the perturbed three-qubit NHH, according to Eq. (1) and Eq. (44) for a perturbation $\epsilon=0.001$. This figure corresponds to the unperturbed case shown in Fig. 4(d). The upper (lower) row of the panels shows the scenario when the dynamical winding is performed in the counterclockwise $\omega$ (clockwise $\omega$) direction. The time period is chosen as $T=500$ [arb. units], , the remaining system parameters are the same as in Fig. 4.
Appendix G Numerical simulations of state dynamics and solution stiffness
In this section, we address the issue of solution stiffness in the simulated state dynamics. Specifically, we observe that when calculating the fidelity for the state-switching between $\psi_{7}$ and $\psi_{8}$, as shown in Fig. 4, the fidelity might exhibit rapid oscillations for certain encircling loops. For instance, in Fig. 13, we plot the state fidelity for the same scenario as in Fig. 4, but with the encircling loop along a circle of radius $r=1$. One can see that during the transition $\psi_{7}\to\psi_{8}$, after crossing the branch cut, the fidelity undergoes rapid oscillations with a non-negligible amplitude, which gradually vanishes as the system approaches the final state $\psi_{8}$.
This stiffness issue arises because the energy sheets $E_{7}$ and $E_{8}$ are very smooth and lie very close to each other when the evolving states transition between these levels near the branch cut. Consequently, the simulated state dynamics may exhibit rapid variations, which can be suppressed by taking very small integration steps when solving the Schrödinger equation, or by varying the shape of the enclosing loop such that the stiffness can be mitigated (as shown in the main text). The interested reader may refer to Ref. [87], where the problem of stiff equations is discussed in greater detail.
In Fig. 14, we plot the energy levels as a function of $\alpha$, corresponding to the scenario shown in Fig. 4. As seen in panel (a), for the first six eigenstates, the energy dependence around the branch cut $\alpha=0$ is linear. This linear dependence ensures stable dynamics, as confirmed in Fig. 13. However, for the states $\psi_{7}$ and $\psi_{8}$, the energy curves are significantly smoother near $\alpha=0$, as shown in Fig. 14(b). Indeed, by expanding the energies $E_{7,8}$ in a power series in $\alpha$, they take a general form $E_{7,8}=u_{2}^{7,8}\alpha^{2}+u_{3}^{7,8}\alpha^{3}+\dots$, where $u^{7,8}_{i}\in{\mathbb{R}}$. This smoothness, i.e., the nearly zero slope of the energy curve, and proximity of the energy levels can thus lead to rapid oscillations in the solutions for the state dynamics. However, as we have demonstrated in the main text, this numerical difficulty can be overcome by choosing, e.g., elliptic loops in the state trajectories, characterized by the large major semi-axis along the $Oy$ axis, which increases the relative slope between the energy curves near the EP and, therefore, can restore the solution stability (see also Fig. 15 and its comparison with Figs. 4 and 13).
Figure 13: Fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between the evolving state $|\psi(t)\rangle$ and the initial eigenstate $|\psi_{k}\rangle$, $k=1,\dots,8$, of the NHH. This figure corresponds to the Fig. 4 in the main text, with only difference that the chosen trajectory loop is a circle with the radius $r=1$ in Eq. (IV.1). The rest system parameters are the same. For trajectory the dynamics for the states $|\psi_{7}\rangle$ and $|\psi_{8}\rangle$ exhibits rapid oscillations in the time interval $T/2<t<T$, due to stiffness of the solution (see the text in the Appendix G for more details).
Figure 14: Energy levels of the three-qubit system, studied in Sec. IV in the main text, as a function of the parameter $\alpha$. (a) The energy levels, corresponding to the first six eigenstates $|\psi_{i}\rangle$, $i=1,\dots,6$, are linear in the vicinity of the branch cuts $\alpha=0$ and have large slopes with respect to zero, ensuring the solution stability of the state dynamics. (b) The energy levels of the eigenstates $|\psi_{7,8}\rangle$ have mostly zero slope near the branch cut and lie close to each other. This can result in rapid oscillations in the state dynamics for some chosen trajectories, as shown in Fig. 13 (see also the text in the Appendix G for more details).
Figure 15: Energy levels corresponding to the eigenstates $|\psi_{7}\rangle$ (black curves) and $|\psi_{8}\rangle$ (brown curves) along the two different loops around the EP in the system parameter space, according to Eq. (IV.1). The energy levels along the elliptic loop, shown as brown solid and black dash-dotted curves, correspond to the scenario in Fig. 4. The energy levels along the circle with the radius $r=1$, shown as brown dotted and black dotted curves, correspond to the scenario in Fig. 13. For the circle trajectory, the energy levels are very smooth and lie close to each other which can lead to rapid oscillations in the state dynamics (as shown in Fig. 13). These oscillations can be mitigated by choosing an appropriate trajectory, e.g., an elliptic loop, where the relative slope between the two energy levels becomes larger.